OpenStudy (anonymous):

two bodies are thrown vertically upwards from walls of height 2.2 m and 1.05m with an intial velocity of 10 m/s ratio of their time of flight is (g=10m/s)

OpenStudy (anonymous):

You have to find two times... up and down, total distance before it strikes the ground? or time in the air? or before it stops mid-air? t1 and t2 knowns- vi- initial velocity= 10 m/s g= 9.8 m/s^2 velocity final = 0 m/s d1= 2.2 m ; d2= 1.05 Is any of this helping?

OpenStudy (anonymous):

Stop mid air is also, referred to as maximum height... so what are they asking for?

OpenStudy (anonymous):

max height is at Vf= o so h; then use the law of motions.

OpenStudy (anonymous):

ok thanxx